1. A bus initially with a speed of 25 m/s slows down with an acceleration of -3 m/s2. How far does it travel before coming to a complete stop?
1
Expert's answer
2021-10-26T17:18:47-0400

Solution.

"v=25 m\/s;"

"a=-3 m\/s^2;"

"S -?;"

"S=\dfrac{v^2-v_0^2}{2a};"

"S=\dfrac{0-625}{-6}=104.17m;"

Answer: "S=104.17m."